1, You plan to purchase a house for $183,000 using a 15-year mortgage obtained from your local bank. You will make a down payment of 10 percent of the purchase price. You will not pay off the mortgage early. Assume the homeowner will remain in the house for the full term and ignore taxes in your analysis.

a, Your bank offers you the following two options for payment. Which option should you choose?

Option 1: Mortgage rate of 5.2 percent and zero points.
Option 2: Mortgage rate of 5 percent and 2 points.

b, Your bank offers you the following two options for payment. Which option should you choose?

Option 1: Mortgage rate of 5.1 percent and 1 point.
Option 2: Mortgage rate of 4.93 percent and 1.5 points.
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Loan Required = Purchase Price*(1-down payment)

Loan Required = 183000*(1-10%) = 164700

a)

Option 1

Loan to be taken = Loan Required /(1-basis point)

Loan to be taken = 1647000/(1-0%)

Loan to be taken = 164700

Monthly rate = 5.2%/12

Monthly Payment = pmt(rate,nper,pv,fv)

Monthly Payment = pmt(5.2%/12,360,-164700,0)

Monthly Payment = $ 904.39

Option 2

Loan to be taken = Loan Required /(1-basis point)

Loan to be taken = 164700/(1-3%)

Loan to be taken = 169793.81

Monthly rate = 5%/12

Monthly Payment = pmt(rate,nper,pv,fv)

Monthly Payment = pmt(5%/12,360,-169793.81,0)

Monthly Payment = $ 911.49

Answer

Option 1 Should be choosen as its monthly payment is lower

b)

Option 1

Loan to be taken = Loan Required /(1-basis point)

Loan to be taken = 164700/(1-1%)

Loan to be taken = 166363.63

Monthly rate = 5.1%/12

Monthly Payment = pmt(rate,nper,pv,fv)

Monthly Payment = pmt(5.1%/12,360,-166363.63,0)

Monthly Payment = $ 903.27

Option 2

Loan to be taken = Loan Required /(1-basis point)

Loan to be taken = 164700/(1-1.5%)

Loan to be taken = 167208.12

Monthly rate = 6.9%/12

Monthly Payment = pmt(rate,nper,pv,fv)

Monthly Payment = pmt(4.93%/12,360,-167208.12,0)

Monthly Payment = $ 890.47

Answer

Option 2 Should be choosen as its monthly payment is lower